If an object is projected vertically upward from an altitude of s0 feet with an initial velocity of v0 ft/sec, then its distance above the ground after t seconds is given by the function below. If s(2) = 136 and s(5) = 55, what are v0 and s0? (Round your answer to the nearest whole number.)
Answer by Alan3354(69443)

No function is shown.
If an object is projected vertically upward from an altitude of s0 feet with an initial velocity of v0 ft/sec, then its distance above the ground after t seconds is given by the function below. If s(2) = 136 and s(5) = 55, what are v0 and s0? (Round your answer to the nearest whole number.)
-------------
I'll use:
s(t) = -16t^2 + vt + s0
s(2) = -64 + 2v + s0 = 136
s(5) = -400 + 5v + s0 = 55
-------------
2v + s0 = 200
5v + s0 = 455
------------- Subtract
-3v = -255
v = 85 ft/sec
----------
s0 = 30 feet